5. A 1.00 L sample of dry air at 250,c and 786 mm Hg contains 0.925 g N2 plus other gases including oxygen, argon, and carbon dioxide. A. What is the partial pressure (in mm Hg) of N2 in the air sample? B. What is the mole fraction and mole percent of N2 in the mixture? (Daltons law of partial pressures) 6. Calculate the rms speed of O2 molecules in a tank at 21°C and 15.7 at
Answer #1

using ideal gas equation ....we have

PV =nRT

n =PV/RT

n = (786/760) X 1.00 / 0.0821 X 298

= 0.04227

and given that moles of N2 present = 0.925/ 28

= 0.0330

hence excess moles ar ethe mole of other gases present ..that is ..0.04227 - 0.0330

= 0.00923 moles

or moles fraction of N2 = moles of N2 / total moles

= 0.0330/ 0.04227

= 0.7806

or acc to daltons partial pressure law

pN2 = XN2 X Ptotal

= 0.7806 X 786

= 613.62 mm of Hg

2)

rms =( 3RT/M)0.5

R = 8.314 ; T = 273+ 21 = 294 k

M = 0.032 kg

hence we have

(3X 8.314 X 294 / 0.032)0.5

= 478.7 m/s